# Zero-shot denoising via neural compression: Theoretical and algorithmic framework

### v1

@@ 6.3 Proof of Theorem 2
Given the input signal ${\bm{x}}\in\operatorname{\mathbb{R}}^{n}$ and ${\bf c}\in\mathcal{C}$, let $\mathrm{Poisson}(\alpha x_{i})$ and $\mathrm{Poisson}(\alpha c_{i})$ denote the distributions corresponding to independent Poisson random variables with respective means $\alpha x_{i}$ and $\alpha c_{i}$. Note that with the chain rule for relative entropy we have
@@
$D_{\text{KL}}(\mathrm{Poisson}(\alpha{\bm{x}})\|\mathrm{Poisson}(\alpha\hat{\bm{x}}))$ $\leq D_{\text{KL}}(\mathrm{Poisson}(\alpha{\bm{x}})\|\mathrm{Poisson}(\alpha\tilde{\bm{x}}))+t_{1}+t_{2},$ (45)
@@
To finish the proof, we bound ($\mathbb{P}\left\lparen\mathcal{E}_{1}\cap\mathcal{E}_{2}\right\rparen$ ) and set $t_{1}$ and $t_{2}$.
To bound ($\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\right\rparen$ ), we apply Lemma 2, where for each ${\bf c}$, we set $w_{i}({\bf c})=\log{1\over c_{i}}$. Then,
@@
($\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\right\rparen$ ) $\leq 2^{R}\exp(-\frac{t_{1}^{2}}{2(\alpha\beta^{2}\sum_{i=1}^{n}x_{i}+\beta t_{1}/3)})$
@@
To bound ($\mathbb{P}\left\lparen\mathcal{E}_{2}^{c}\right\rparen$ ), we again apply Lemma 2, with $w_{i}=\log{1\over{\tilde{x}}_{i}}$, and derive
($\mathbb{P}\left\lparen\mathcal{E}_{2}^{c}\right\rparen$ ) $\leq\exp(-\frac{t_{2}^{2}}{2(\alpha\beta^{2}\sum_{i=1}^{n}x_{i}+\beta t_{2}/3)})$
@@
($\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\cap\mathcal{E}_{2}^{c}\right\rparen\leq 2^{R}\exp(-\frac{t_{1}^{2}}{4n\alpha\beta^{2}})+\exp(-\frac{t_{2}^{2}}{4n\alpha\beta^{2}}).$ ) . (49)
@@
($\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\cap\mathcal{E}_{2}^{c}\right\rparen\leq 2^{-\eta R}.$ (50)
6.4 Proof of Theorem 3
@@
Following the similar setup as in Section 6.2, we get
@@
($\mathbb{P}\left\lparen\left\lvert\sum^{n}_{i=1}(y_{i}-\alpha x_{i})e_{i}^{({\bf c})}\right\rvert\geq t\right\rparen\leq 2\exp\left\lparen-\frac{t^{2}}{2(\sigma_{n}^{2}+w_{M}t/3)}\right\rparen,$ (55)
@@
($\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\right\rparen$ ) $\leq 2^{R+1}\exp\left(-\frac{t_{1}^{2}}{2(nx_{\max}^{3}\alpha+t_{1}x_{\max}/3)}\right)$ (57)
@@
($\mathbb{P}\left\lparen\mathcal{E}_{2}^{c}\right\rparen\leq 2\exp\left(-\frac{t_{2}^{2}}{2(nx_{\max}^{3}\alpha+t_{2}x_{\max}/3)}\right).$ ) . (58)
@@
($\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\cap\mathcal{E}_{2}^{c}\right\rparen\leq 2^{R+1}\exp(-\frac{t_{1}^{2}}{4n\alpha})+2\exp(-\frac{t_{2}^{2}}{4n\alpha}).$ ) . (59)
@@
($\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\cup\mathcal{E}_{2}^{c}\right\rparen\leq 2^{-\eta R+2}.$ (60)
@@
$\leq\|{\bm{d}}\|_{2}^{2}+\sqrt{{nR\over\alpha}}\left\lparen 4\sqrt{\ln 2}\right\rparen\left\lparen\sqrt{1+\eta}+\sqrt{\eta}+1\right\rparen.$ ) .
6.5 Proof of Corollary 1
Proof.
First we need to design a lossy compression for the set of signals in $\mathcal{Q}_{n}$, defined as
$\mathcal{Q}_{n}=\left\{{\bm{x}}\in\operatorname{\mathbb{R}}^{n}:\lVert{\bm{x}}\rVert_{0}\leq k,\lVert{\bm{x}}\rVert_{2}\leq 1\right\}.$ (61)
For a $k$-sparse ${\bm{x}}\in\mathcal{Q}_{n}$, let ${\bm{x}}^{(k)}\in\operatorname{\mathbb{R}}^{k}$ denote the $k$-dimensional vector derived from the non-zero coordinates of ${\bm{x}}$. Let $\mathcal{B}^{k}_{2}=\left\{{\bm{a}}\in\operatorname{\mathbb{R}}^{k}:\|{\bm{a}}\|_{2}\leq 1\right\}$, then ${\bm{x}}^{(k)}\in\mathcal{B}^{k}_{2}$. We define a lossy compression code with rate $R$ and distortion $\delta$ as follows:
  1. 1.
    To encode the number of non-zero entries and their locations in ${\bm{x}}\in\mathcal{Q}_{n}$, one spends at most $\log_{2}k+k\log_{2}n$ bits.
  2. 2.
    To encode the values of the non-zero coordinates, the number of balls of radius $\varepsilon$ required to cover $\mathcal{B}^{k}_{2}$, $N(\mathcal{B}^{k}_{2},\varepsilon)$ is bounded as
@@ 6.5 Proof of Corollary 1
    This implies that at most by using
    $k\log_{2}\frac{1}{\varepsilon}\leq\log_{2}N(\mathcal{B}^{k}_{2},\varepsilon)\leq k\log_{2}\left\lparen\frac{2}{\varepsilon}+1\right\rparen$ ) (62)
    bits, we can encode the values of the non-zero entries of ${\bm{x}}\in\mathcal{Q}_{n}$, or equivalently all the entries of ${\bm{x}}^{(k)}\in\mathcal{B}^{k}_{2}$, with error less than $\varepsilon$.
Note, as defined in Section 3, we measure the distortion of a lossy compression code as squared error. Therefore, to achieve the defined distortion $\delta$, we need to set $\varepsilon=\sqrt{\delta}$, which implies that the overall rate is bounded as
$\log_{2}k+k\log_{2}n+k\log_{2}\left\lparen\frac{1}{\sqrt{\delta}}\right\rparen\leq R\leq\log_{2}k+k\log_{2}n+k\log_{2}\left\lparen\frac{2}{\sqrt{\delta}}+1\right\rparen,$ (63)
using the upper bound on the rate as in (63) in combination with Theorem 1, we have
${1\over\sqrt{n}}\|{\bm{x}}-\hat{\bm{x}}\|_{2}$ $\leq\sqrt{\delta}+2\sigma_{z}\sqrt{(2\ln 2)\left\lparen\log_{2}k+k\log_{2}n+k\log_{2}\left\lparen\frac{2}{\sqrt{\delta}}+1\right\rparen\right\rparen\over n}(1+2\sqrt{\eta})$ )
$\leq\sqrt{\delta}+2\sigma_{z}\sqrt{(2\ln 2)\left\lparen\log_{2}k+k\log_{2}n+{k\over 2}\log_{2}({1\over\delta})+k\log_{2}(2+\sqrt{\delta})\over n\right\rparen}(1+2\sqrt{\eta}),$
setting $\delta={1\over n}$, with probability at least $1-2^{-\eta R+2}$ we have
$\frac{1}{\sqrt{n}}\lVert\hat{{\bm{x}}}-{\bm{x}}\rVert$ $\leq\frac{1}{\sqrt{n}}+\sigma_{z}C\sqrt{\log_{2}k+{k\over 2}\log_{2}n+k\log_{2}(2+1/\sqrt{n})\over n}$
$=\frac{1}{\sqrt{n}}+\sigma_{z}C\sqrt{\frac{k\log_{2}n}{n}}\sqrt{1+\frac{2\log_{2}k}{k\log_{2}n}+\frac{2\log_{2}(2+1/\sqrt{n})}{\log_{2}n}}$
$\leq\frac{1}{\sqrt{n}}+\sigma_{z}C\sqrt{\frac{k\log_{2}n}{n}}\left\lparen 1+\frac{\log_{2}k}{k\log_{2}n}+\frac{\log_{2}(2+1/\sqrt{n})}{\log_{2}n}\right\rparen$ )
$=\sigma_{z}C\sqrt{\frac{k\log_{2}n}{n}}+\left\lparen\frac{1}{\sqrt{n}}+\frac{\sigma_{z}C\log_{2}k}{2\sqrt{n}}+\frac{\sigma_{z}C\sqrt{k}\log_{2}(2+1/\sqrt{n})}{\sqrt{n}}\right\rparen,$
where $C=2\sqrt{\ln 2}(1+2\sqrt{\eta})$. Squaring both sides of the last inequality gives
$\frac{1}{n}\lVert\hat{{\bm{x}}}-{\bm{x}}\rVert^{2}\leq\sigma_{z}^{2}C^{2}\frac{k\log_{2}n}{n}+o(1),$ (64)
which holds with probability at least $1-\frac{4}{\left\lparen kn^{3k/2}\right\rparen^{\eta}}$, using the lower bound on the rate in (63). ∎
7 Conclusions
We have studied maximum likelihood compression-based denoising, and provided theoretical characterization of its performance under both AWGN and Poisson noise. Furthermore, we introcued ZS-NCD, a new zero-shot neural-compression-based denoising and demonstrated that it achieves state-of-the-art performance among zero-shot methods, in both AWGN and Poisson denoising.
Acknowledgment
A.Z., X.C., S.J. were supported by NSF CCF-2237538.
Appendix A Appendix
@@ A.1 Network structure
For our experiments we used 3 convolutional layers in the encoder with 128 number of channels for the first two layers in the encoder (and the last two layers of decoder), see Figure 5. For color images we choose the number of channels in the last encoder (and first decoder) layer equals to 32, and for grayscale images equals to 16. The MLP-based network of the ablation study in Section 5 has 3 fully connected layers in the encoder with 1024 hidden units for the first two layers in the encoder (and the last two layers of decoder). The number of hidden units in the last encoder (and first decoder) layer equals to 16. As activation function we use GDN [29] for Conv network and ReLU for MLP.
A.2 Additional experiments on Poisson denoising
Unknown noise level
In a more realistic case, we generally have no access to the noise level $\alpha$ in Poisson noise model, which makes the normalization using true $\alpha$ not practical. We propose to use the estimated $\hat{\lambda}$ for normalization both in MSE and MLE based denoising. Given that $\mathbb{E}\left[y_{i}\right]=\alpha x_{i}$, so within a n-dimensional single noisy image, $\frac{1}{n}\sum^{n}_{i=1}y_{i}\approx\alpha\frac{1}{n}\sum^{n}_{i=1}x_{i}$, where in most of the cases, we assume $\mathbb{E}\left[{\bm{x}}\right]\approx 0.5$, so $\hat{\alpha}\approx 2\frac{1}{n}\sum^{n}_{i=1}y_{i}$. We then use the estimated noise level $\hat{\alpha}$ to normalize both MSE and MLE based optimization for denoising Poisson noise. See Table 4 for the result of this noise parameter estimation on a sample image.
Setting the hyperparameter $\lambda$ under Poisson noise
Having an estimate of noise parameter $\alpha$, we can decompose the MSE between ${\bm{y}}$ and the denoised image $\hat{\bm{x}}$ as
$={1\over n\alpha}\sum_{i=1}^{n}x_{i}+{1\over n}\|{\bm{x}}-\hat{\bm{x}}\|_{2}^{2}+{2\over n}\mathbb{E}\left[({\bm{y}}-\alpha{\bm{x}})^{T}({\bm{x}}-\hat{\bm{x}})\right],$ (65)
with same logic described in Section 4, assuming that the second and third terms in (65) are close to zero for a close estimate $\hat{\bm{x}}$ to ${\bm{x}}$, then ${1\over n}\mathbb{E}\left[\|{\bm{y}}/\alpha-\hat{\bm{x}}\|_{2}^{2}\right]\approx\frac{1}{2\alpha}$ with the assumption that $\mathbb{E}\left[{\bm{x}}\right]\approx 0.5$ (as used in $\alpha$ estimation), we can use $\|\hat{\bm{x}}^{(k)}-\bm{y}\|^{2}>\frac{1}{2\alpha}$ as the decision rule in the Algorithm 1 to set the hyperparameter $\lambda$. We have empirically observed that MSE(${\bm{y}}/\alpha$,$\hat{\bm{x}}$) in training our networks is close to $\frac{1}{2\alpha}$ as reported in Table 4, which indicates that $\frac{1}{2\alpha}$ is a good approximation of the MSE that can be used as threshold in selecting $\lambda$. When $\alpha$ is not known, we obtain the estimate $\hat{\alpha}$, and use $\frac{1}{2\hat{\alpha}}$ as a valid threshold to search $\lambda$.
true $\alpha$ estimated $\hat{\alpha}$ empirical MSE(${\bm{y}}/\alpha$,$\hat{\bm{x}}$) [dB] $1/(2\alpha)$ [dB] $1/(2\hat{\alpha})$ [dB]
25 23.02 17.12 16.98 16.63
50 46.05 19.66 20.00 19.64
@@ A.2 Additional experiments on Poisson denoising
MSE and likelihood estimation without knowing true $\alpha$
@@
We compare the MSE and MLE distortion for Poisson denoising using the described estimated $\hat{\alpha}$ in Table 5.
@@
A.3 Study on factors in patch-wise compression affecting denoising
In this section, we explain the intuition behind why learning compression networks and denoising on overlapped patches is feasible. The centered pixels in the patches are better compressed as empirically observed in [55], thus they can provide better denoising performance. To study the contribution of each patch containing the single pixel to be denoised we design the experiment that, in the denoising phase, we denoise the overlapped patches, but only a single pixel at the same location from each patch is used to construct the final denoised image, instead of averaging all of them as in (8). We show the denoising performance of each pixel location in Figure 6. The PSNR at each pixel denotes the denoising performance of only using the specific pixel of each overlapped patches with stride 1. We can find that the boundary pixels give lower PSNR, which is consistent with previous research findings that the centered pixels are better compressed. Next, we analyze the effect of patch size in both learning and denoising phases. Given that scaler quantization is applied and the entropy model is learned on latent code of the patches, the compression performance on the latent code is affected by both the patch size and the number of downsampling operations in CNN-based encoder. We design the experiment that 3 downsampling operators are applied to patch size 8 and 16, where the latent code sizes are $1\times 1\times n_{b}$ and $2\times 2\times n_{b}$ respectively, where the denoising performance at each pixel location is in Figure 6 (Left) and (Middle), and if we increase the downsampling to 4 for patch size 16, which results in the latent code size to be $1\times 1\times n_{b}$, the denoising performance is in Figure 6 (Right). We find that spatial size of the latent code to be quantized matters given the scaler quantization limitation, the reconstructed output by the decoder will be restricted by the only correlated latent code as we can observe. Motivated from this, we perform the learning and denoising phases both patch-wise with proper networks structure, all pixels in each patch are used and the overlapped areas are averaged properly to reduce the variance of the compression-based estimates.
@@ A.3 Study on factors in patch-wise compression affecting denoising
Appendix B Additional numerical results
@@ B.3 Kodak24 Dataset
B.4 Microscopy Mouse Nuclei Dataset
For these images with noise level $\sigma_{z}=(10,20)$ we set $\lambda=(200,600)$, we train the networks for 20K steps to obtain the results. We report the detailed denoising performance in Table 12 and 13 respectively.
Table 12: Denoising performance under AWGN $\mathcal{N}(0,\sigma^{2}I)$ on fluorescence microscopy dataset: Mouse Nucle. Images are cropped into $128\times 128$. Noise level $\sigma_{z}=10$.
# JPEG2K BM3D DIP DD ZS-N2N ZS-N2S S2S ZS-NCD

### v2

@@ A.4 Proof of Theorem 2
Given the input signal ${\bm{x}}\in\mathbb{R}^{n}$ and ${\bf c}\in\mathcal{C}$, let $\mathrm{Poisson}(\alpha{\bm{x}})$ and $\mathrm{Poisson}(\alpha{\bf c})$ denote the distributions corresponding to independent Poisson random variables with respective means $\alpha x_{i}$ and $\alpha c_{i}$. Note that
@@
$)\displaystyle D_{\text{KL}}(\mathrm{Poisson}(\alpha{\bm{x}})\|\mathrm{Poisson}(\alpha\hat{\bm{x}}$)) $\leq D_{\text{KL}}(\mathrm{Poisson}(\alpha{\bm{x}})\|\mathrm{Poisson}(\alpha\tilde{\bm{x}}))+t_{1}+t_{2},$ (48)
@@
[…]
@@
[…]
@@
[…]
@@
($).\displaystyle\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\cap\mathcal{E}_{2}^{c}\right\rparen\leq 2^{R}\exp(-\frac{t_{1}^{2}}{4n\alpha\beta^{2}})+\exp(-\frac{t_{2}^{2}}{4n\alpha\beta^{2}}$). (52)
@@
[…]
A.5 Proof of Theorem 3
@@
Following the similar setup as in Section A.2, we get
@@
[…]
@@
[…]
@@
($).\displaystyle\mathbb{P}\left\lparen\mathcal{E}_{2}^{c}\right\rparen\leq 2\exp\left(-\frac{t_{2}^{2}}{2(nx_{\max}^{3}\alpha+t_{2}x_{\max}/3)}\right$). (61)
@@
($).\displaystyle\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\cap\mathcal{E}_{2}^{c}\right\rparen\leq 2^{R+1}\exp(-\frac{t_{1}^{2}}{4n\alpha})+2\exp(-\frac{t_{2}^{2}}{4n\alpha}$). (62)
@@
[…]
@@
Appendix B Additional experiments and experimental settings
@@ B.1 Network structure
For our experiments we used 3 convolutional layers in the encoder with 128 number of channels for the first two layers in the encoder (and the last two layers of decoder), see Figure 5. For color images we choose the number of channels in the last encoder (and first decoder) layer equals to 32, and for grayscale images equals to 16. The MLP-based network of the ablation study in Section 5 has 3 fully connected layers in the encoder with 1024 hidden units for the first two layers in the encoder (and the last two layers of decoder). The number of hidden units in the last encoder (and first decoder) layer equals to 16. As activation function we use GDN [balle2017endtoend] for Conv network and ReLU for MLP.
B.2 More recent improvements on Zero-shot Noise2Noise
Zero-shot Noise2Noise (ZS-N2N) [mansour2023zsn2n] learns a mapping between two downsampled noisy images extracted from the original image. This strategy of generating noisy samples was extended in two recent concurrent works. Dual-sampling Noise2Noise (DS-N2N) [bai2025dual] identifies a key sub-optimality in ZS-N2N: the network learns to denoise at a lower resolution but is then directly applied to the larger, original resolution. To address this, the authors proposed to apply an additional bicubic upsampling to the downsampled images and train the network on both the low-resolution and the upsampled pairs, which boosts the performance of ZS-N2N.
In Pixel2Pixel [ma2025pixel2pixel], the authors identified the local sampling in ZS-N2N insufficient to break the spatial correlation of real-world noise. To address this issue, Pixel2Pixel first finds a “pixel bank” for each pixel based on the non-local similar patches. Then, by randomly choosing pixels from the bank, it generates multiple noisy pairs to be used as a pseudo-training dataset for zero-shot denoising.
We have compared the performance of both methods under AWGN and Poisson noise in Tables 5 and 6, respectively. Each paper’s official code was used to report the numbers in the tables. In AWGN, our compression-based denoiser ZS-NCD outperforms both methods while achieving the second-best performance in Poisson denoising. Pixel2Pixel performs well in Poisson denoising.
Table 5: Average PSNR(dB)/SSIM denoising performance on Kodak24 dataset under AWGN $\mathcal{N}(0,\sigma_{z}^{2})$. Best results are in bold, second-best are underlined.
$\sigma_{z}=15$ $\sigma_{z}=25$ $\sigma_{z}=50$
ZS-N2N (2023) 32.30 / 0.8650 29.54 / 0.7798 25.82 / 0.6151
DS-N2N (2025) 32.31 / 0.8803 29.64 / 0.8044 25.42 / 0.6378
Pixel2Pixel (2025) 31.31 / 0.8707 29.89 / 0.8098 26.55 / 0.6873
ZS-NCD 33.18 / 0.9026 30.60 / 0.8144 27.89 / 0.7464
Table 6: Average PSNR(dB)/SSIM denoising performance on Kodak24 dataset under Poison noise, $\mathrm{Poisson}(\alpha{\bm{x}})/\alpha$. Best results are in bold, second-best are underlined.
$\alpha=15$ $\alpha=25$ $\alpha=50$
ZS-N2N (2023) 26.80 / 0.6757 28.21 / 0.7374 30.13 / 0.8076
DS-N2N (2025) 27.29 / 0.7016 28.50 / 0.7540 30.27 / 0.8250
Pixel2Pixel (2025) 28.22 / 0.7390 29.26 / 0.7891 30.41 / 0.8372
ZS-NCD 27.64 / 0.7432 28.77 / 0.7677 30.60 / 0.8235
@@ B.2 More recent improvements on Zero-shot Noise2Noise
B.3 MSE and likelihood estimation under Poisson noise without knowing true $\bm{\alpha}$
@@
We compare the MSE and MLE distortion for Poisson denoising in Table 7 using the estimated $\hat{\alpha}$ as explained in Section 4.
@@
B.4 Study on factors in patch-wise compression affecting denoising
In this section, we explain the intuition behind why learning compression networks and denoising on overlapped patches is feasible. The centered pixels in the patches are better compressed as empirically observed in [he2021checkerboard], thus they can provide better denoising performance. To study the contribution of each patch containing the single pixel to be denoised we design the experiment that, in the denoising phase, we denoise the overlapped patches, but only a single pixel at the same location from each patch is used to construct the final denoised image, instead of averaging all of them as in (9). We show the denoising performance of each pixel location in Figure 6. The PSNR at each pixel denotes the denoising performance of only using the specific pixel of each overlapped patches with stride 1. We can find that the boundary pixels give lower PSNR, which is consistent with previous research findings that the centered pixels are better compressed. Next, we analyze the effect of patch size in both learning and denoising phases. Given that scaler quantization is applied and the entropy model is learned on latent code of the patches, the compression performance on the latent code is affected by both the patch size and the number of downsampling operations in CNN-based encoder. We design the experiment that 3 downsampling operators are applied to patch size 8 and 16, where the latent code sizes are $1\times 1\times n_{b}$ and $2\times 2\times n_{b}$ respectively, where the denoising performance at each pixel location is in Figure 6 (Left) and (Middle), and if we increase the downsampling to 4 for patch size 16, which results in the latent code size to be $1\times 1\times n_{b}$, the denoising performance is in Figure 6 (Right). We find that spatial size of the latent code to be quantized matters given the scaler quantization limitation, the reconstructed output by the decoder will be restricted by the only correlated latent code as we can observe. Motivated from this, we perform the learning and denoising phases both patch-wise with proper networks structure, all pixels in each patch are used and the overlapped areas are averaged properly to reduce the variance of the compression-based estimates.
@@ B.4 Study on factors in patch-wise compression affecting denoising
Appendix C Additional numerical results
@@ C.3 Kodak24 Dataset
C.4 Microscopy Mouse Nuclei Dataset
For these images with noise level $\sigma_{z}=(10,20)$ we set $\lambda=(200,600)$, we train the networks for 20K steps to obtain the results. We report the detailed denoising performance in Table 14 and 15 respectively.
Table 14: Denoising performance under AWGN $\mathcal{N}(0,\sigma^{2}I)$ on fluorescence microscopy dataset: Mouse Nuclei. Images are cropped into $128\times 128$. Noise level $\sigma_{z}=10$.
# JPEG2K BM3D DIP $D$ ZS-N2N ZS-N2S S2S ZS-NCD